OpenStudy (anonymous):

Which of the following is not equivalent to the formula for the circumference of a circle, C = 2 π r ?

OpenStudy (primeralph):

Yeah?

OpenStudy (primeralph):

medal?

OpenStudy (anonymous):

options are c/rπ=2 c/2π=r c/2r=π 2π/c=r

OpenStudy (primeralph):

what do you think?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

The last one isn't

OpenStudy (anonymous):

@hea u think its the last one?

OpenStudy (primeralph):

@hea I'm going to have to report you.

OpenStudy (anonymous):

why?

OpenStudy (primeralph):

not allowed to post answers

OpenStudy (anonymous):

Oh yeah, sorry

OpenStudy (primeralph):

@JULIANCHICO I don't make the rules. And if you have to insult me for obeying the rules, you will get suspended.

OpenStudy (primeralph):

@.Sam. please settle this

OpenStudy (primeralph):

@amistre64

OpenStudy (primeralph):

Why did you delete the comment?| You lack guts.